A series R L circuit is connected to a 110V, 50 Hz ac source. If the voltage across the resistor is 85 V, 

find the power dissipated in the circuit, if the value of inductor is 0.05 H


1
Expert's answer
2021-05-30T14:39:30-0400

Let us introduce the notation

"U=110 V,f=50 Hz, U_R=85 V, L=0.05H"

1).Determine the resistance of the inductor

"X_L=2 \cdot \pi \cdot f \cdot L=2 \cdot 3.14 \cdot 50 \cdot 0.05=15.71 \Omega"

2).Determine the voltage across the inductor

"U_L=\sqrt{U^2-U_R^{2}}=\sqrt{110^2-85^{2}}=\sqrt{12100-7225}=69.82V"

3).Determine the current in the circuit

"I=\frac{U_L}{X_L}=\frac{69.82}{15.71}=4.45A"

4).Determine the total power of the circuit

"S=U \cdot I=110 \cdot4.45=489.5VA"

5).Determine the reactive power of the circuit

"Q=I^2 \cdot X_L=4.45^2 \cdot 15.71=311.1var"

5).Determine the active power dissipated in the circuit

"P=\sqrt{S^2-Q^{2}}=\sqrt{489.5^2-311.1^{2}}=377.93W"